TB Problem Qu. 3-252 Baab Corporation is a ...

Baab Corporation is a manufacturing firm that uses job-order costing. The company's inventory balances were as follows at the beginning and end of the year:

Beginning Balance Ending Balance
Raw materials $ 14,950 $ 22,950
Work in process $ 27,950 $ 9,950
Finished Goods $ 62,950 $ 77,950
  • The company applies overhead to jobs using a predetermined overhead rate based on machine-hours. At the beginning of the year, the company estimated that it would work 33,950 machine-hours and incur $302,155 in manufacturing overhead cost. The following transactions were recorded for the year:
  • Raw materials were purchased, $315,950.
  • Raw materials were requisitioned for use in production, $307,950 ($280,050 direct and $27,900 indirect).
  • The following employee costs were incurred: direct labor, $377,950; indirect labor, $96,950; and administrative salaries, $172,950.
  • Selling costs, $147,950.
  • Factory utility costs, $10,950.
  • Depreciation for the year was $184,000 of which $173,000 is related to factory operations and $11,000 is related to selling, general, and administrative activities.
  • Manufacturing overhead was applied to jobs. The actual level of activity for the year was 34,190 machine-hours.
  • Sales for the year totaled $1,325,000.

Required:

a. Prepare a schedule of cost of goods manufactured.

b. Was the overhead underapplied or overapplied? By how much?

c. Prepare an income statement for the year. The company closes any underapplied or overapplied overhead to Cost of Goods Sold.
Answer #1

Part A

Baab Company

Schedule of Cost of Goods Manufactured

Direct materials

Raw materials inventory, beginning

14950

Add: Purchases of raw materials

315950

Total raw materials available

330900

Deduct: Raw materials inventory, ending

22950

Raw materials used in production

307950

Deduct: Indirect materials

27900

Direct materials

280050

Direct labor

377950

Manufacturing overhead applied (302155/33950*34190)

304291

Total manufacturing costs

962291

Add: Beginning work in process inventory

27950

990241

Deduct: Ending work in process inventory

9950

Cost of goods manufactured

980291

Part B

underapplied   overhead

$4509

Actual overhead = 27900+96950+10950+173000 = 308800

underapplied   overhead = 308800-304291

Part C

Baab Company

Income Statement

Sales

1325000

Cost of goods sold

969800

Gross margin

355200

Selling and administrative expenses:

Administrative salaries

172950

Selling costs

147950

Depreciation

11000

331900

Net operating income

$23300

Cost of goods sold = 62950+980291-77950+4509 = 969800